A myopic person has no astigmatism, but has a near-point of 15 cm and a far point of 50 cm. Can you prescribe a single pair of glasses that will give him normal vision? this is supposed to be Ans: (-2 diopter-lens)?

in the solution I checked, their was a solution for both far and near points, but their is only one answers, so shouldn't there be only one solution
Answer #1

Read your question carefully...A myopic person has no astigmatism, but has a near-point of 15 cm and a far point of 50 cm. Can you prescribe a single pair of glasses that will give him normal vision?

That's why you need to calculate for two lenses.

That's what I think Buddy!!

Using lens equation: 1/f = 1/q + 1/p

For image at 25 cm and since it will be on the same side of the lens as the image at 15.

Apply 1/f = 1/p + 1/q

1/f = 1/(15) + 1/(-25)

f = 37.5 cm

Thus, power = 1/.375 = +2.668 Diopters

For the far point,

we should have images at infinity and form them at 50 cm

1/f = 0 + 1/(-50)

f = -50 cm

The dipotic power is 1/-.5 = -2.00 Diopters.

Thus, the person needs bifocals with powers of +2.67 and -2.00.